The Pen, Evan, and Torves Partnership has asked you to assist in winding-up its business affairs. You compile the following information:

  1. The partnership’s trial balance on June 30, 20X1, is
Debit Credit
Cash $ 5,800
Accounts Receivable (net) 21,500
Inventory 11,000
Plant and Equipment (net) 97,900
Accounts Payable $ 18,100
Pen, Capital 53,100
Evan, Capital 43,000
Torves, Capital 22,000
Total $ 136,200 $ 136,200
  1. The partners share profits and losses as follows: Pen, 60 percent; Evan, 20 percent; and Torves, 20 percent.
  2. The partners are considering an offer of $101,000 for the firm’s accounts receivable, inventory, and plant and equipment as of June 30. The $101,000 will be paid to creditors and the partners in installments, the number and amounts of which are to be negotiated.


Required:
Prepare a cash distribution plan as of June 30, 20X1, showing how much cash each partner will receive if the partners accept the offer to sell the assets.
Answer #1

Answer :

Cash Distribution Plan
June 30, 20X1
Loss Absorption Power Capital Accounts
Pen   Evan Torves Pen   Evan Torves
Profit and Loss percentages 60% 20% 20%
Preliquidation capital balances       (53,100)       (43,000)       (22,000)
Loss Absorption Power (Capital balances / Loss percent) **      (88,500)      (215,000)      (110,000)
Decrease highest LAP to next highest :
Evan 215,000 to 110,000 = 105,000*0.20       105,000        21,000
     (88,500)      (110,000)      (110,000)       (53,100)       (22,000)       (22,000)
Decrease highest LAP to next highest :
Evan    - 110,000 to 88,500 = 21,500 * 0.2         21,500          4,300
Torves - 110,000 to 88,500 = 21,500 * 0.2         21,500          4,300
     (88,500)       (88,500)       (88,500)       (53,100)       (17,700)       (17,700)
Summary of Cash Distribution
(If Offer of $101,000 is accepted)
Accounts Payable Pen Evan Torves
Cash Available (101,000+5,800) $               106,800
First                   (18,100)                     18,100
Next                   (21,000)         21,000
Next                     (8,600)           4,300 4300
Additional in Profit & loss ratio(6:2:2)                   (59,100)         35,460 11820 11820
                          -   $                 18,100 $     35,460 $     37,120 $     16,120

** Working of Loss absorption power:

Loss Absorption Power:
Pen 53,100/0.6         88,500
Evan 43,000/0.2       215,000
Torves 22,000/0.2       110,000